mikhailjrr:

Milla and Luka are 3 kilometers apart and walking toward each other. Milla's average speed is 5 kilometers per hour and Luka's is 4 kilometers per hour. A table showing Rate in kilometers per hour, Time in hours, and Distance in kilometers. The first row shows, Milla, and has 5, t, and 5 t. The second row shows, Luka, and has 4, t, and 4 t.

Mercury:

Which equation can be used to find t, the time it takes for Milla and Luka to meet? 5t + 4t = 1 5t + 4t = 3 5t – 4t = 0 5t – 4t = 3

Mercury:

if somebody travels (x) km/hr in (y) hrs, they'll cover xy km applying this logic, Milla travels 5t km in t hours, and Luka travels 4t together, they must cover 3km in order for them to meet therefore, 5t + 4t = 3